# Self-supervised pretraining of vision transformers for animal behavioral analysis and neural encoding

### v1

@@ 1 Introduction
Understanding the relationship between brain and behavior is a fundamental challenge across a wide range of medical and scientific disciplines [25, 10, 41]. Precise methods for extracting meaningful information from behavioral videos are essential for advancing these fields [49]. Self-supervised learning has revolutionized image and video understanding through large-scale foundation models [9, 7, 53, 15], offering powerful tools that are beginning to transform scientific analyses [19, 31, 83]. However, these models have yet to be effectively translated to specialized domains like animal behavior analysis, creating a significant opportunity for methods that bridge cutting-edge machine learning with the specific demands of neuroscience and behavioral research.
@@
Animal behavior videos present unique characteristics and challenges distinct from general video understanding. Controlled experiments generate large quantities of videos with static backgrounds and consistent camera angles, where the primary variation arises from animal movements and interactions. Behavioral videos enable numerous downstream analyses, and here we focus on three fundamentally different applications that collectively address a large proportion of behavioral neuroscience use cases: (1) neural encoding, which requires extracting behavioral features that correlate with simultaneously recorded brain activity [10, 49, 66]; (2) pose estimation, which tracks specific anatomical landmarks for quantitative analysis of movement patterns [41, 49]; and (3) action segmentation, which classifies distinct behavioral states like grooming, rearing, or social interactions on every frame [10, 49]. Each task demands different representations of the same underlying behavioral data, yet current approaches typically require task-specific models and extensive labeled datasets.
The primary barrier to sophisticated behavioral analysis in most laboratories is not the conceptual challenge of creating experiment-specific datasets, but rather the practical constraints of scale and technical expertise [67]. Individual labs typically lack sufficient data volume to train high-quality models from scratch and face computational limitations that make large-scale model development infeasible [49, 67]. This creates substantial adoption barriers across the scientific community. While recent approaches attempt to reduce labeled data requirements [40, 48, 57, 60, 3, 14, 63] or eliminate them entirely through unsupervised methods [2, 18, 61, 39, 73, 68], each comes with significant limitations. Supervised methods still require tedious and potentially biased human annotation, while unsupervised approaches often require complex post-processing to yield interpretable results suitable for scientific inquiry. Both approaches are typically task-specific and therefore lack versatility.
@@
Our work addresses these challenges through a novel self-supervised pretraining framework for vision transformers that serves as a robust backbone for multiple downstream tasks. Beast (BEhavioral Analysis via Self-supervised pretraining of Transformers) leverages the unique properties of experimental videos by combining masked autoencoding [15] to capture rich frame-level appearance information with temporal contrastive learning [9] to model behavioral dynamics. We introduce a novel frame sampling strategy for the contrastive loss, specifically designed to focus on learning representations of animal behavior against static experimental backgrounds. Beast trains on videos from a single experimental setup, creating tailored, versatile models that can be fine-tuned for multiple analytical needs specific to that experimental context. We demonstrate the value of this approach through comprehensive evaluation on three downstream tasks: (1) neural activity prediction (or “neural encoding”) in three mouse datasets; (2) pose estimation across four datasets spanning two species and single- and multi-camera setups; and (3) action segmentation in both single- and multi-animal experiments. Beast achieves competitive or superior performance across these tasks while eliminating the need for pose estimation or other preprocessing steps typically required by existing neural encoding and action segmentation methods. These results establish Beast as a simple yet powerful foundation that can accelerate behavioral understanding across disciplines where fine-grained analysis is essential. Code is available at https://github.com/paninski-lab/beast.
@@ 2 Related Work
Neural encoding measures how observable signals predict neural activity, and provide a quantitative framework for interrogating neural representations. Earlier approaches applied generalized linear models to single neurons using controlled stimuli such as visual or auditory inputs [46, 65, 50, 42]. More recently, deep learning methods have shown great promise in predicting neural population responses to sensory stimuli, including visual [76, 56, 69], auditory [23, 35], and tactile [84] inputs. The widespread adoption of video monitoring during experiments has demonstrated that video-based behavioral covariates explain significant neural variability in both spontaneous [58, 63] and task-driven behaviors [44, 29, 71, 8, 82]. For example, Musall et al. [44] showed that uninstructed movements explain a substantial fraction of cortical population variance, and the International Brain Lab’s Brainwide Map [29] leveraged large-scale, region-resolved encoding analyses to chart the distribution of task-related information across the brain. However, extracting rich spatiotemporal information from video remains challenging. Most studies rely on either a small set of keypoints [63, 29, 71, 8] or latent dimensions using PCA [58, 44] or autoencoders [1, 71, 8], with limited efforts to predict neural activity directly from raw video [71].
@@
Large-scale models for animal behavior analysis have predominantly focused on single tasks. For pose estimation, methods differ in how they balance flexibility and labeling requirements. DeepLabCut [40] leverages ImageNet-pretrained backbones for fine-tuning on experiment-specific labeled datasets, offering flexibility but requiring more manual labeling. In contrast, several specialized pose estimation tools provide tailored solutions for common experimental setups, such as top-down views of freely moving mice (SuperAnimal [77]) and facial analysis of head-fixed rodents (Facemap [63]), significantly reducing labeling requirements. Similarly, in action segmentation, specialized systems developed for resident-intruder assays [57, 14] achieve high performance but remain limited to a specific experimental paradigm. While VideoPrism [83] offers a general foundation model supporting multiple behavioral tasks [62], it relies on a frozen backbone trained on generic internet data rather than domain-specific content, and remains unavailable to the public. Despite these advances, a critical gap remains: there are currently no accessible solutions for creating general behavior analysis models that leverage abundant unlabeled data to improve performance across multiple tasks. Our work, Beast, addresses this need by introducing a pretraining strategy that enables individual labs to develop experiment-specific models supporting diverse downstream tasks using their own unlabeled videos. This approach allows researchers to extract maximum value from their experimental data without extensive manual labeling.
@@
Contrastive learning has emerged as a powerful self-supervised representation learning framework; among its many predecessors and variants, SimCLR [9] popularized a simple and effective recipe that maximizes agreement between differently augmented views of the same sample via a contrastive loss in latent space. The contrastive method has also been extended to the video domain [52, 54]. Masked modeling frameworks are a complementary approach that have demonstrated remarkable success, particularly masked autoencoding (MAE) [15], which revolutionized visual self-supervised learning by adapting BERT-style masked prediction to images using Vision Transformers [12]. This approach has been extended to video data by VideoMAE [64] and BEVT [70], which leverage spatiotemporal dependencies. Various works have combined contrastive and MAE objectives as a more efficient alternative for capturing spatiotemporal dependencies [43, 20, 38, 17, 34]. Of note is ViC-MAE [17], which uses local (patch-based) features for a masked image modeling loss. The local features are also pooled into a global feature vector which is used with a contrastive loss computed across frames from multiple videos. We further describe this method in the next section.
@@ 3 Methods
Beast uses a combination of image-based masked autoencoding (MAE) and contrastive objectives (Fig. 1A). The MAE loss excels at capturing per-frame appearance details, while the temporal contrastive loss captures dependencies across frames. This integration enables a single backbone to excel across diverse downstream tasks, from precise keypoint localization to predicting complex structure in neural activity (Fig. 1B).
Beast builds upon ViC-MAE [17], which combines masked autoencoding and contrastive losses, but introduces key adaptations for neuroscience applications. The most significant modification is how frames are sampled for the contrastive loss. ViC-MAE allows any two frames from the same video to be a positive pair, and frames from different videos are negative pairs [75]. While this may be appropriate for benchmark datasets with short clips, animal behavior experiments generate continuous, long-duration recordings where behaviors repeat across time. Our approach instead defines positive frames within a narrow temporal window around the anchor (±$\pm 1$ frame), while allowing negative frames to be either distant and dissimilar frames in the same video, or from entirely different videos. Crucially, we find this strategy outperforms that of ViC-MAE (Table 5). See Appendix B for more details on our frame selection strategy and additional training and architecture simplifications of ViC-MAE.
Figure 1: Beast framework. A: Our self-supervised pretraining framework Beast combines masked autoencoding [15] with temporal contrastive learning [9]. An anchor frame at time t is paired with a positive frame from $t\pm 1$, while more distant frames from the same video, or frames from other videos, serve as negative examples. Frames are divided into patches, with most patches randomly masked. A Vision Transformer (ViT) processes the remaining patches, which must reconstruct all patches. The ViT’s CLS tokens, which serve as a global representation of each frame, are nonlinearly projected into a new space where the contrastive loss pulls anchor-positive pairs together while pushing anchor-negative pairs apart. B: Beast supports various downstream neuro-behavioral tasks including neural encoding, pose estimation, and action segmentation.
@@
The standard image ViT [12] data pipeline starts with a 2D image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ (where $H$, $W$, $C$ are height, width, channels) and splits it into 2D patches, each with shape $(P\times P\times C)$, where the patch size $P$ is typically 16. Each patch is reshaped to a vector of length $P^{2}C$, and all patches are concatenated into a sequence of the $N$ flattened 2D patches $\mathbf{x}_{p}\in\mathbb{R}^{N\times(P^{2}C)}$, where $N$ is the total number of patches. Each flattened patch is mapped with a trainable linear projection to a “patch token,” a vector of size $D=768$ (the standard embedding size for the ViT-B/16 architecture). We add standard 1D position embeddings to the patch tokens to retain information about the patch location. We prepend a learnable CLS token to the sequence of patch tokens, which serves as a summary representation for the image. The resulting patch tokens augmented with position embeddings ($\mathbf{t}\in\mathbb{R}^{N\times D}$) and the concatenated CLS token serve as the input to the standard ViT encoder architecture.
@@
The masked autoencoding loss randomly masks out a high proportion of the patch tokens (here, 0.75 [15]). We call the resulting unmasked tokens $\mathbf{t}_{um}\in\mathbb{R}^{L\times D}$, where $L=0.25\times N$ is the number of unmasked tokens. The unmasked tokens are then processed by the ViT, resulting in embeddings $\mathbf{z}_{um}=\textsc{ViT}(\mathbf{t}_{um})$. The masked embeddings $\mathbf{t}_{m}\in\mathbf{R}^{(N-L)\times D}$ (consisting of all zeros) are then combined with the unmasked embeddings processed by ViT to reform a complete patch sequence $\mathbf{z}\in\mathbb{R}^{N\times D}$. These embeddings are passed through a transformer decoder to produce a reconstruction $\hat{\mathbf{x}}_{p}\in\mathbb{R}^{N\times P^{2}C}$ trained via mean square error (MSE): $\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{p=1}^{N}(\mathbf{x}_{p}-\hat{\mathbf{x}_{p}})^{2}$. We refer to the model trained only with this masked autoencoding loss as ViT-M.
@@
The masked autoencoding loss is sufficient for reconstructing low-level features on individual frames. To imbue our embeddings with temporal information (which may be required for certain downstream tasks), we employ a contrastive loss that produces similar embeddings for frames close in time, and distinct embeddings for frames far apart in time or from different videos. To achieve this, each batch with $B$ samples contains $B/2$ anchor frames, and each anchor frame $\mathbf{x}_{t}^{v}$ (from time $t$ and video $v$) has a randomly chosen positive frame from $\mathbf{x}_{t\pm 1}^{v}$. All remaining $B-2$ frames are treated as negative frames. Note this approach differs from other temporal contrastive losses that allow any frame from the anchor frame’s video to be a positive frame [75, 17], which does not perform well with temporally-extended behavioral videos (Table 5). To improve the robustness of this approach, we select the initial set of anchor frames from a given video to be as visually distinct from each other as possible (Table 4). We utilize the InfoNCE loss [45] computed on nonlinear projections of the CLS embeddings that are output by the ViT. The projector outputs $\{\mathbf{z}^{p}_{b}\}$ are used for the contrastive learning, calculated as $\mathcal{L}_{\text{InfoNCE}}=-\frac{2}{B}\sum_{i\in\mathcal{A}}\log\frac{\exp(\mathbf{z}_{i}^{p}\cdot\mathbf{z}_{i^{\prime}}^{p})}{\sum_{j\neq i}\exp(\mathbf{z}_{i}^{p}\cdot\mathbf{z}_{j}^{p})}$, where $i^{\prime}$ is the positive example associated with $i$ and $\mathcal{A}$ is the set of $B/2$ anchor frames. We refer to the model trained with both the masked autoencoding and contrastive losses as Beast.
@@
We initialize our models with pretrained ImageNet weights [11, 15]. Details of the training dataset, data augmentations, and batch construction are provided in Appendix B. We define the loss as $\mathcal{L}_{\text{MSE}}+\lambda\cdot\mathcal{L}_{\text{InfoNCE}}$, where $\lambda$ is a hyperparameter that balances the two losses and is set between $0.01$ and $0.03$. We select $\lambda$ using the validation sets of the various datasets. Models are trained for 800 epochs using the AdamW optimizer [37] with a cosine annealing learning rate scheduler [36], taking approximately 25 hours on 8 Nvidia A40 GPUs.

### v2

# arXiv:2507.09513v2[q-bio.NC] 27 Feb 2026

@@ 1 Introduction
Understanding the relationship between brain and behavior is a fundamental challenge across a wide range of medical and scientific disciplines [Krakauer et al., 2017, Datta et al., 2019]. Precise methods for extracting meaningful information from behavioral videos are essential for advancing these fields [Pereira et al., 2020]. Self-supervised learning has revolutionized image and video understanding through large-scale foundation models [Chen et al., 2020, Caron et al., 2021, He et al., 2022], offering powerful tools that are beginning to transform scientific analyses [Huang et al., 2023a, Lastufka et al., 2024]. However, these models have yet to be effectively translated to specialized domains like animal behavior analysis, creating a significant opportunity for methods that bridge cutting-edge machine learning with the specific demands of neuroscience and behavioral research.
@@
Animal behavior videos present unique characteristics and challenges distinct from general video understanding. Controlled experiments generate large quantities of videos with static backgrounds and consistent camera angles, where the primary variation arises from animal movements and interactions. These videos enable numerous downstream analyses, and here we focus on three fundamentally different applications that collectively address a large proportion of behavioral neuroscience use cases: (1) neural activity prediction (or “neural encoding”), which requires extracting behavioral features that correlate with simultaneously recorded brain activity [Datta et al., 2019, Pereira et al., 2020, Urai et al., 2022]; (2) pose estimation, which tracks specific anatomical landmarks for quantitative analysis of movement patterns [Mathis and Mathis, 2020, Pereira et al., 2020]; and (3) action segmentation, which classifies distinct behavioral states like grooming, rearing or social interactions on every frame [Datta et al., 2019, Pereira et al., 2020]. Each task demands different representations of the same underlying behavioral data, and current approaches typically require task-specific models and extensive labeled datasets [von Ziegler et al., 2021]. Furthermore, most approaches fail to leverage the vast amounts of unlabeled data generated by behavior experiments, a significant untapped resource that, if harnessed properly, could substantially improve performance on these downstream tasks.
@@
We address these challenges through a novel self-supervised pretraining framework for raw videos that produces a robust backbone for multiple downstream neuro-behavioral tasks. Beast (BEhavioral Analysis via Self-supervised pretraining of Transformers) leverages the unique properties of experimental videos by combining masked autoencoding [He et al., 2022] to capture rich frame-level appearance information with temporal contrastive learning [Hyvarinen and Morioka, 2016] to model behavioral dynamics. We introduce a novel frame sampling strategy for the contrastive loss, designed to focus on learning representations of animal behavior against static backgrounds. Beast trains on videos from a single experimental setup, creating tailored, versatile models that can be fine-tuned for multiple analytical needs specific to that experimental context. We demonstrate the value of this approach through comprehensive evaluation on three downstream tasks: (1) neural encoding in three mouse datasets; (2) pose estimation across four datasets spanning two species and single- and multi-view setups; and (3) action segmentation in both single- and multi-animal setups. Beast achieves competitive or superior performance for the neural encoding and action segmentation tasks–outperforming baselines such as DINOv2 [Oquab et al., 2023]–while eliminating the need for the pose estimation step typically required by existing methods, dramatically reducing manual labeling effort. At the same time, pose estimation remains valuable for producing interpretable features critical to understanding movement dynamics, and Beast enables significantly improved pose estimation for a given labeling budget. These results establish Beast as a simple yet powerful foundation that can accelerate behavioral understanding across disciplines where fine-grained analysis is essential.
@@ 2 Related Work
Neural encoding measures how observable signals predict neural activity, and provide a quantitative framework for interrogating neural representations. Earlier approaches applied generalized linear models to single neurons using controlled stimuli such as visual or auditory inputs [Paninski, 2004, Truccolo et al., 2005, Pillow et al., 2008, McFarland et al., 2013]. More recently, deep learning methods have shown great promise in predicting neural population responses to sensory stimuli, including visual [Yamins et al., 2014, Schrimpf et al., 2018, Wang et al., 2025a], auditory [Kell et al., 2018, Li et al., 2023], and tactile [Zhuang et al., 2017] inputs. The widespread adoption of video monitoring during experiments has demonstrated that video-based behavioral covariates explain significant neural variability in both spontaneous [Stringer et al., 2019, Syeda et al., 2024] and task-driven behaviors [Musall et al., 2019, IBL et al., 2025a, Wang et al., 2025b, Chen et al., 2024, Zhang et al., 2025b]. For example, Musall et al. [2019] showed that uninstructed movements explain a substantial fraction of cortical neural variance, and the International Brain Lab leveraged large-scale, region-resolved encoding analyses to chart the distribution of task-related information across the brain [IBL et al., 2025a]. However, extracting rich spatiotemporal information from video remains challenging. Most studies rely on either a small set of keypoints [Syeda et al., 2024, IBL et al., 2025a, Wang et al., 2025b, Chen et al., 2024] or latent dimensions using PCA [Stringer et al., 2019, Musall et al., 2019] or autoencoders [Batty et al., 2019, Wang et al., 2025b, Chen et al., 2024], with limited efforts to predict neural activity directly from raw video (but see Wang et al. [2025b]).
@@
Large-scale models for animal behavior analysis have predominantly focused on single tasks. For pose estimation, methods differ in how they balance flexibility and labeling requirements. DeepLabCut [Mathis et al., 2018] leverages ImageNet-pretrained backbones for fine-tuning on experiment-specific labeled datasets, offering flexibility but requiring more manual labeling. This work inspired a range of other general-purpose animal pose estimation tools including LEAP [Pereira et al., 2019], DeepPoseKit [Graving et al., 2019], TRex [Walter and Couzin, 2021], SLEAP [Pereira et al., 2022], and Lightning Pose [Biderman et al., 2024]. In contrast, several specialized pose estimation tools provide tailored solutions for common experimental setups, such as top-down views of freely moving mice [Ye et al., 2024] and facial analysis of head-fixed rodents  [Syeda et al., 2024, Daruwalla et al., 2024], significantly reducing labeling requirements. Similarly, in action segmentation, specialized systems developed for resident-intruder assays [Segalin et al., 2021, Goodwin et al., 2024] achieve high performance but remain limited to a specific experimental paradigm. While VideoPrism [Zhao et al., 2024] offers a general foundation model supporting multiple behavioral tasks [Sun et al., 2024], it relies on a frozen backbone trained on generic internet data rather than domain-specific content. Despite these advances, no accessible solutions exist for creating general behavior analysis models that leverage unlabeled data across multiple tasks. Beast addresses this gap by enabling individual labs to develop experiment-specific models from their own unlabeled videos for diverse analyses.
@@
Contrastive learning has emerged as a powerful self-supervised learning (SSL) framework; among its many predecessors and variants, SimCLR [Chen et al., 2020] popularized a simple and effective recipe that maximizes agreement between differently augmented views of the same sample via a contrastive loss in latent space. The contrastive method has also been extended to the temporal [Hyvarinen and Morioka, 2016] and video domains [Qian et al., 2021, Recasens et al., 2021, Dave et al., 2022]. Another line of self-supervised approaches uses knowledge distillation, where a student network learns to match the outputs of a teacher network, such as DINO [Caron et al., 2021, Oquab et al., 2023, Siméoni et al., 2025]. Masked modeling is a complementary approach that has demonstrated remarkable success, particularly masked autoencoding (MAE) [He et al., 2022], which revolutionized visual self-supervised learning by adapting BERT-style masked prediction to images using Vision Transformers [Dosovitskiy et al., 2020]. VideoMAE [Tong et al., 2022] and BEVT [Wang et al., 2022] extended this approach to video data by leveraging spatiotemporal dependencies. Various works have combined contrastive and MAE objectives as a more efficient alternative for capturing spatiotemporal dependencies, as video models can require much more compute for training and inference [Mishra et al., 2022, Huang et al., 2023b, Lu et al., 2023, Lehner et al., 2024]. Of note is ViC-MAE [Hernandez et al., 2024], which uses patch-based features for a masked autoencoding loss. The local features are also pooled into a global feature vector which is used with a contrastive loss computed across frames from multiple videos. The efficiency of contrastive-based methods compared to native video models is of particular interest in our application domain, where labs often do not have access to extensive compute resources. SSL techniques from computer vision have increasingly been adapted to the study of animal behavior. These approaches use SSL to extract useful feature representations from pose, image, or video data, which are then applied to downstream tasks such as action segmentation (Appendix F).
@@ 3 Methods
Beast uses a combination of an image-based masked autoencoding (MAE) loss–which excels at capturing per-frame appearance details–and temporal contrastive loss–which captures dependencies across frames (Fig. 1A). This integration enables a single backbone to excel across diverse downstream tasks, from precise keypoint localization to predicting complex structure in neural activity (Fig. 1B).
Beast builds upon ViC-MAE [Hernandez et al., 2024], which combines masked autoencoding and contrastive losses, but introduces key adaptations for neuroscience applications. The most significant modification is how frames are sampled for the contrastive loss. ViC-MAE allows any two frames from the same video to be a positive pair, and frames from different videos are negative pairs [Xu and Wang, 2021]. While this may be appropriate for benchmark datasets with short clips, animal behavior experiments generate long-duration recordings where behaviors repeat across time. We instead define positive frames within a narrow temporal window around the anchor ($\pm 1$ frame), while allowing negative frames to be either distant and dissimilar frames in the same video, or from different videos. Crucially, this strategy outperforms that of ViC-MAE (Table 5). See Appendix B for more details on our frame selection strategy and additional training and architecture simplifications of ViC-MAE.
Figure 1: Beast framework. A: Our self-supervised pretraining framework Beast combines masked autoencoding [He et al., 2022] with temporal contrastive learning [Chen et al., 2020]. An anchor frame at time t is paired with a positive frame from $t\pm 1$, while more distant frames from the same video, or frames from other videos, serve as negative examples. Frames are divided into patches, with most patches randomly masked. A vision transformer (ViT) processes the remaining patches, which must reconstruct all patches. The ViT CLS tokens, which serve as a global representation of each frame, are nonlinearly projected to a new space where the contrastive loss pulls anchor-positive pairs together and pushes anchor-negative pairs apart. B: Beast supports various downstream neuro-behavioral tasks including neural encoding, pose estimation, and action segmentation.
@@
The standard image ViT [Dosovitskiy et al., 2020] data pipeline starts with a 2D image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ ($H$, $W$, $C$ are height, width, channels) and splits it into 2D patches, each with shape $(P\times P\times C)$, where the patch size $P$ is typically 16. Each patch is reshaped to a vector of length $P^{2}C$, and all patches are concatenated into a sequence of the $N$ flattened 2D patches $\mathbf{x}_{p}\in\mathbb{R}^{N\times(P^{2}C)}$. Each flattened patch is mapped with a trainable linear projection to a “patch token,” a vector of size $D$. We add 1D position embeddings to the patch tokens to retain patch location information. We add a learnable CLS token to the patch token sequence, which serves as a global representation for the image. The resulting patch tokens augmented with position embeddings ($\mathbf{t}\in\mathbb{R}^{N\times D}$) and the concatenated CLS token serve as the input to the standard ViT encoder.
@@
The masked autoencoding (MAE) loss randomly masks out a high proportion of the patch tokens (here, 0.75 [He et al., 2022]). We call the resulting unmasked tokens $\mathbf{t}_{um}\in\mathbb{R}^{L\times D}$, where $L=0.25\times N$ is the number of unmasked tokens. The unmasked tokens are processed by the ViT to produce embeddings $\mathbf{z}_{um}=\textsc{ViT}(\mathbf{t}_{um})$. The masked embeddings $\mathbf{z}_{m}\in\mathbf{R}^{(N-L)\times D}$ (consisting of all zeros) are then combined with the unmasked embeddings processed by ViT to form a complete patch sequence $\mathbf{z}\in\mathbb{R}^{N\times D}$, which is passed through a transformer decoder to produce a reconstruction $\hat{\mathbf{x}}_{p}\in\mathbb{R}^{N\times P^{2}C}$ trained via mean square error: $\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{p=1}^{N}(\mathbf{x}_{p}-\hat{\mathbf{x}_{p}})^{2}$. We refer to the model trained only with this MAE loss as ViT-M.
@@
The masked autoencoding loss is sufficient for reconstructing low-level features on individual frames. To imbue our embeddings with temporal information (which may be required for certain downstream tasks), we employ a contrastive loss that produces similar embeddings for frames close in time, and distinct embeddings for frames far apart in time or from different videos. To achieve this, each batch with $B$ samples contains $B/2$ anchor frames, and each anchor frame $\mathbf{x}_{t}^{v}$ (from time $t$ and video $v$) has a randomly chosen positive frame from $\mathbf{x}_{t\pm 1}^{v}$. All remaining B2B-2 frames are treated as negative frames. Note this approach differs from other temporal contrastive losses that allow any frame from the anchor frame’s video to be a positive frame [Xu and Wang, 2021, Hernandez et al., 2024], which does not perform well with temporally-extended behavioral videos (Table 5). To improve the robustness of this approach, we select the initial set of anchor frames from a given video to be as visually distinct from each other as possible (Table 4). We utilize the InfoNCE loss [Oord et al., 2018] computed on nonlinear projections of the CLS embeddings (which outperform other frame aggregation methods, Table 7) that are output by the ViT. The projector outputs $\{\mathbf{z}^{p}_{b}\}$ are used for the contrastive learning, calculated as $\mathcal{L}_{\text{InfoNCE}}=-\frac{2}{B}\sum_{i\in\mathcal{A}}\log\frac{\exp(\mathbf{z}_{i}^{p}\cdot\mathbf{z}_{i^{\prime}}^{p})}{\sum_{j\neq i}\exp(\mathbf{z}_{i}^{p}\cdot\mathbf{z}_{j}^{p})}$, where $i^{\prime}$ is the positive example associated with $i$ and $\mathcal{A}$ is the set of $B/2$ anchor frames. We refer to the model trained with both the masked autoencoding and contrastive losses as Beast.
@@
We initialize our models with pretrained ImageNet weights [Deng et al., 2009, He et al., 2022]. Details of dataset construction, data augmentations, and batch construction are provided in Appendix B. We define the loss as $\mathcal{L}_{\text{MSE}}+\lambda\cdot\mathcal{L}_{\text{InfoNCE}}$, where $\lambda$ balances the two losses and is selected using the validation sets of the various datasets. Models are trained for 800 epochs using the AdamW optimizer [Loshchilov and Hutter, 2017] with a cosine annealing learning rate scheduler [Loshchilov and Hutter, 2016], taking approximately 25 hours on 8 Nvidia A40 GPUs.